# A Resource Efficient Quantum Kernel

### v3

@@ Abstract
Quantum processors may enhance machine learning by mapping high-dimensional data onto quantum systems for processing. Conventional feature maps, for encoding data onto a quantum circuit are currently impractical, as the number of entangling gates scales quadratically with the dimension of the dataset and the number of qubits. In this work, we introduce a quantum feature map designed to handle high-dimensional data with a significantly reduced number of qubits and entangling operations. Our approach preserves essential data characteristics while promoting computational efficiency, as evidenced by extensive experiments on benchmark datasets that demonstrate a marked improvement in both accuracy and resource utilization when using our feature map as a kernel for characterization, as compared to state-of-the-art quantum feature maps. Our noisy simulation results, combined with lower resource requirements, highlight our map’s ability to function within the constraints of noisy intermediate-scale quantum devices. Through numerical simulations and small-scale implementation on a superconducting circuit quantum computing platform, we demonstrate that our scheme performs on par or better than a set of classical algorithms for classification. While quantum kernels are typically stymied by exponential concentration, our approach is affected with a slower rate with respect to both the number of qubits and features, which allows practical applications to remain within reach. Our findings herald a promising avenue for the practical implementation of quantum machine learning algorithms on near future quantum computing platforms.
@@ I Introduction
High-dimensional data is prevalent in modern machine learning tasks, including image and speech recognition, natural language processing, and medical diagnostics. While classical machine learning techniques can handle these high-dimensional problems, they often require substantial computational resources, particularly as the dimensionality of the data increases [1, 2, 3]. Quantum computing has emerged as a promising avenue to address these challenges. Quantum kernel methods, for instance, have shown potential for accelerating data analysis by efficiently learning relationships in high-dimensional spaces encoded into quantum states [4, 5, 6, 7, 8, 9]. To harness large-dimensional Hilbert spaces for processing data, quantum computers need to efficiently encode classical data onto quantum states. These quantum feature maps are required for quantum support vector classification [10, 11, 12, 13, 14] and quantum neural networks [10, 11, 15, 16, 14], including data reuploading [17], and are germane to essentially all quantum machine learning paradigms. Yet, the most popular feature maps require numbers of qubits and controlled-not (CNOT) gates linear and quadratic in the number of data features, respectively, severely shortening their practical application. We introduce a feature map whose qubit requirement is reduced by a factor of at least two and whose CNOT-gate requirement is reduced to linear in the number of qubits, and show these improvements to be sufficient for practical application on current devices.
@@
Here, after introducing our quantum feature map with the reduced resource requirements, we examine its performance using a kernel method in classification of various datasets. We compare our results with several classical methods and the commonly used Pauli feature map known as the ZZ feature map and consistently show competing performance. We present a detailed background of quantum feature maps and their challenges in Section I.1. In Section II.1 we introduce our approach. We lay out our experimental methodology and results in Section II.2, including small-scale tests on IBM’s ibm_quebec and IBM_torino machines. Our results stimulate considerations of quantum machine learning as an early practical application of application-starved NISQ devices. In our work, we utilize this feature map within a predefined quantum kernel function known as a the fidelity quantum kernel for the support vector classifier (SVC) algorithm in Qiskit [21, 14]; as a result, we will refer to both the feature map and the kernel as CPMap.
@@ I.1.2 Pauli Feature Maps
and enable the map to encode both individual contributions and pairwise interactions between features into an entangled quantum state. Owing to its capacity to enrich the expressivity of the quantum feature space, the ZZFeatureMap is one of the most commonly used data encoding methods in quantum classification tasks [23, 24, 25, 22, 4].
@@ I.1.3 Challenges and Limitations
Figure 2: CPMap vs. ZZfeatureMap based on quantum resources. (a) Circuit depth vs. number of features (b) Number of CNOTs used in the circuit vs. number of features (c) Number of qubits required vs. number of features for both of the features maps for different number of features in the dataset
@@ II.1.1 Resource Complexity
Encoding a data point with $F$ features using the ZZFeatureMap typically requires $n=F$ qubits and $O(F^{2})$ entangling gates. In CPMap, we need $n\approx F/2$ qubits to encode $F$ features, and around O(F)~O(F) CNOTs, a substantial reduction versus ZZ. This resource saving, together with reduced qubit count ( Eq. (15)), is the key reason CPMap is attractive for NISQ devices; see Fig. 2 for a visual comparison.
@@ II.2.1 Results with Seven Features After PCA
Figure 4: MCC-score-based analysis of different kernels on different datasets scaled to each have 7 features, comparing the effectiveness of each kernel in handling specific types of data. The CPMap (second from right for each dataset) always outperforms the ZZFeatureMap( rightmost for each dataset) and always performs better than one or more of the classical kernels. For some datasets, certain kernels completely failed, so there is no bar visible for the ZZFeatureMap for the INS and PD datasets and for the sigmoid kernel for the BCD dataset. Dataset acronyms: Ionosphere (INS), Breast Cancer Diagnostic (BCD), Credit Card Fraud ($C$), Parkinson’s Disease (PD), Stellar Classification (SC), Heart Disease (HD), and Titanic Survival (T).
For some datasets, we repeated the CPMap up to two times but uploaded the same features for each repetition. This increases the total number of CNOT gates by a factor of two but maintains the total number of qubits and the significant reduction in the number of CNOT gates; see Appendix C.1 for details.
@@ II.2.2 Results for the Noisy Simulation with Seven Features
As demonstrated in Figure 5, our CPMap consistently surpassed the noiseless performance of ZZFeatureMap, even in the presence of noise.
@@ II.2.3 Results for the Noisy Simulation with sixteen Features
II.2.4 A case for quantum advantage
@@
The consistent outperformance by CPMap across diverse datasets suggests a promising avenue for achieving quantum advantage in predictive tasks. The substantial margin by which CPMap leads for datasets with large numbers of features hints at its unique ability to capture complex patterns that classical kernels might not discern as effectively, thereby bolstering the hypothesis that quantum machine learning could offer computational benefits over traditional algorithms.
@@ II.3 Statistical Significance Testing
In addition to demonstrating quantum advantage, we further validated the performance of our custom CPMap kernel against traditional RBF, poly, and linear kernels using Parkinson’s disease dataset. We employed statistical significance testing to confirm the superiority of CPMap in terms of accuracy and the MCC.
@@
This statistical analysis and visualization confirm the significant advantage of the CPMap kernel over traditional kernels in the context of Parkinson’s disease data, reinforcing the potential for quantum machine learning to achieve computational superiority in predictive tasks.
@@ II.4.3 Test with PCA reduced dataset on ibm_torino
In this case, we deployed our CPMap kernel circuit on IBM’s ibm_torino, a 133-qubit superconducting quantum processor belonging to the Heron family. Heron-class devices feature improvements in qubit coherence and native two-qubit gate implementations. According to IBM Quantum’s platform, this device has a T1 time of $1.83\times 10^{-4}$ s and T2 time of $1.41\times 10^{-4}$ s.
@@
The observed performance gap between the statevector simulation and the real-device execution can be attributed to decoherence, gate errors, and limited circuit depth. Nevertheless, the successful deployment of CPMap on ibm_torino confirms its feasibility on today’s NISQ devices and underscores the promise of quantum kernels in practical machine learning tasks.
Figure 11: Comparison of CPMap based kernel on the curated Parkinson dataset. The simulated CPMap-based Kernel achieves the highest performance, followed by RBF and quantum hardware (ibm_torino).
II.5 Issue of Exponential Concentration:
@@ III Discussion and Conclusion
Our proposed feature map paves the way for the application of quantum machine learning algorithms to more complex and higher-dimensional data, even with limited qubit resources. The CPMap’s remarkable efficiency in qubit utilization, evident through its requirement of approximately $\frac{N}{2}$ qubits to represent $N$ features, marks a substantial improvement over the ZZFeatureMap. This efficiency not only eases the computational load but also enhances the scalability of quantum models. In our comparative analysis with classical kernels such as the RBF, linear, and polynomial kernels, the CPMap sometimes demonstrated superior performance. Our finding demonstrate the potential for quantum algorithms to outperform their classical counterparts under specific conditions, especially when dealing with high-dimensional data and limited training samples.
@@
The implications of our findings extend beyond the immediate scope of this study. The CPMap’s promising performance in efficiently handling high-dimensional datasets opens new pathways for future work, particularly in applying quantum algorithms to more complex and diverse datasets with applications in health sciences to material discovery. Further exploration into its realization existing noisy quantum computing platforms and future small scale fault-tolerant devices, as opposed to simulations, could offer deeper insights into its practical utility and performance in real-world applications.
Our proposed CPMap is expected to replace the existing Pauli feature maps due its significantly improved performance at reduced resource requirements in terms of the number of qubits and CNOT gates with respect to the number of features. This goes beyond a theoretical promise through our numerical demonstrations and small-scale implementation on a noisy quantum computer. We expect the CPMap protocol to play a crucial role in advancement of quantum machine learning and in developing its early applications on realistic problems with performance that can potentially beat existing classical methods.

### v4

@@ Abstract
Quantum processors may enhance machine learning by mapping high-dimensional data onto quantum systems for processing. Conventional feature maps for encoding data onto a quantum circuit are impractical, as the number of entangling gates scales quadratically with the dimension of the dataset. We introduce a quantum feature map designed to handle high-dimensional data with a significantly reduced number of qubits and entangling operations. Our approach preserves essential data characteristics while promoting computational efficiency, evidenced by extensive experiments on benchmark datasets that demonstrate a marked improvement in both accuracy and resource utilization when using our feature map as a kernel for characterization, as compared to state-of-the-art quantum feature maps. Our noisy simulation results highlight our map’s ability to function within the constraints of noisy intermediate-scale quantum devices. Through numerical simulations and small-scale implementation on a superconducting circuit quantum computing platform, we demonstrate that our scheme performs on par or better than a set of classical algorithms for classification. Our approach empirically delays the onset of exponential concentration relative to existing feature maps under the diagnostics considered here, which can improve the practical operating regime of fidelity kernels on near-term devices. Our results indicate that resource-efficient feature maps can broaden the range of kernel-based QML experiments that are feasible on near-term platforms, and motivate further validation on larger datasets and additional hardware.
@@ I Introduction
High-dimensional data is prevalent in modern machine learning tasks, including image and speech recognition, natural language processing, and medical diagnostics. While classical machine learning techniques can handle these high-dimensional problems, they often require substantial computational resources, particularly as the dimensionality of the data increases [1, 2, 3, 4, 5]. Quantum computing has emerged as a promising avenue to address these challenges. Quantum kernel methods, for instance, have shown potential for accelerating data analysis by efficiently learning relationships in high-dimensional spaces encoded into quantum states [6, 7, 8, 9, 10, 11]. To harness large-dimensional Hilbert spaces for processing data, quantum computers need to efficiently encode classical data onto quantum states. These quantum feature maps are required for quantum support vector classification [12, 13, 14, 15, 16] and quantum neural networks [12, 13, 17, 18, 16], including data reuploading [19], and are germane to essentially all quantum machine learning paradigms. Yet, the most popular feature maps require numbers of qubits and controlled-not (CNOT) gates linear and quadratic in the number of data features, respectively, severely shortening their practical application. We introduce a feature map whose qubit requirement is reduced by a factor of at least two and whose CNOT-gate requirement is reduced to linear in the number of qubits, and show these improvements to be sufficient for practical application on current devices.
@@
Here, after introducing our quantum feature map with the reduced resource requirements, we examine its performance using a kernel method in classification of various datasets. We compare our results with several classical methods and the commonly used Pauli feature map known as the ZZ feature map and consistently show competing performance.We emphasize that ZZFeatureMap is used here as a widely available reference implementation of a Pauli-style entangling feature map, rather than as the only relevant baseline. [6, 26, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] Many alternative embeddings exist (including data re-uploading and other hardware-efficient constructions), and to avoid over-interpreting the comparison as “CPMap vs. ZZ”, we additionally benchmark CPMap against a suite of fixed (non-trained) data re-uploading feature maps in Appendix E. We present a detailed background of quantum feature maps and their challenges in Section I.1. In Section II.1 we introduce our approach. We lay out our experimental methodology and results in Section II.2, including small-scale tests on IBM’s ibm_quebec and IBM_torino machines. Our results stimulate considerations of quantum machine learning as an early practical application of application-starved NISQ devices. In our work, we utilize this feature map within a predefined quantum kernel function known as a the fidelity quantum kernel for the support vector classifier (SVC) algorithm in Qiskit [23, 16]; as a result, we will refer to both the feature map and the kernel as CPMap.
@@ I.1.2 Pauli Feature Maps
and enable the map to encode both individual contributions and pairwise interactions between features into an entangled quantum state. Owing to its capacity to enrich the expressivity of the quantum feature space and to the difficulty of simulating it classically, the ZZFeatureMap is one of the most commonly used data encoding methods in quantum classification tasks [25, 26, 27, 24, 6].
@@ I.1.3 Challenges and Limitations
Figure 2: Resource scaling of CPMap vs. ZZFeatureMap (ideal vs. hardware-compiled). The top row reports ideal//logical circuit resources (no transpilation) as a function of the number of features: (a) circuit depth, (b) two-qubit gates count (CNOT) operations, and (c) qubit count required by the feature map. The bottom row reports the same quantities after transpiling the circuits to the IBM Torino backend (native gate set and coupling constraints), where the native two-qubit operation is CZ; thus subplot (e) reports CZ counts for the compiled circuits. Across both settings, CPMap exhibits substantially reduced two-qubit resources and depth growth compared to ZZFeatureMap, and the qualitative scaling advantage is preserved after device-aware compilation.
@@ II.1.1 Resource Complexity
Encoding a data point with $F$ features using the ZZFeatureMap typically requires $n=F$ qubits and approximately $\mathcal{O}(F^{2})$ two-qubit entangling operations, leading to rapidly increasing depth as $F$ grows. In contrast, CPMap encodes $F$ features using substantially fewer qubits, with $n\approx F/2$ (cf. Eq. (15)), and exhibits an approximately linear growth in entangling cost, i.e., $\mathcal{O}(F)$ two-qubit operations in the ideal (pre-compilation) circuit. Figure 2 reports this comparison in two settings: (top row) the logical circuits prior to transpilation, where the two-qubit count is reported in terms of CX (CNOT) gates, and (bottom row) the same circuits after transpilation to the IBM Torino backend, where the native two-qubit operation is CZ and the compiled two-qubit cost is therefore reported as CZ count. Importantly, the qualitative resource advantage of CPMap (lower depth growth, fewer two-qubit gates, and reduced qubit requirements) is preserved under device-aware compilation, making it particularly attractive for near-term (NISQ) implementations.
For completeness, we provide a hardware-feasibility analysis based on backend-timing circuit durations and device coherence times ($T_{1}$, $T_{2}$) in Appendix H (see Fig. 20). This shows that the compiled gate schedule remains well within coherence across the tested feature dimensions.
@@ II.2.1 Results with Seven Features After PCA
Figure 4: MCC-score-based analysis of different kernels on different datasets scaled to each have 7 features, comparing the effectiveness of each kernel in handling specific types of data. The CPMap (second from right for each dataset) outperforms the ZZFeatureMap (rightmost) on all datasets in this 7-feature benchmark, and is competitive with strong classical kernels (linear, poly, RBF, sigmoid), sometimes exceeding them depending on the dataset. For some datasets, certain kernels completely failed, so there is no bar visible for the ZZFeatureMap for the INS and PD datasets and for the sigmoid kernel for the BCD dataset. Dataset acronyms: Ionosphere (INS), Breast Cancer Diagnostic (BCD), Credit Card Fraud ($C$), Parkinson’s Disease (PD), Stellar Classification (SC), Heart Disease (HD), and Titanic Survival (T).
For some datasets, we repeated the CPMap up to two times but uploaded the same features for each repetition. This increases the total number of CNOT gates by a factor of two but maintains the total number of qubits and the significant reduction in the number of CNOT gates; see Appendix C.1 for details. An additional multi-class experiment on the Stellar Classification dataset (balanced $N=1500$) is reported in Appendix F, and an expanded comparison against fixed data re-uploading feature-map variants (DR1–DR6) is provided in Appendix E.
@@ II.2.2 Results for the Noisy Simulation with Seven Features
As demonstrated in Figure 5, CPMap remains competitive under these noise models and, in this benchmark, achieves higher MCC than ZZFeatureMap even when ZZFeatureMap is evaluated in the ideal (noiseless) setting.
@@ II.2.3 Results for the Noisy Simulation with sixteen Features
II.2.4 A case for quantum usefulness
@@
The consistent outperformance of CPMap in our experiments indicates an empirical advantage for this embedding under the studied settings, particularly when feature counts are larger and circuit resources are constrained. The substantial margin by which CPMap leads for datasets with large numbers of features hints at its unique ability to capture complex patterns that classical kernels might not discern as effectively, thereby bolstering the hypothesis that quantum machine learning could offer computational benefits over traditional algorithms.
@@ II.3 Statistical Significance Testing
In addition to demonstrating quantum usefulness, we further validated the performance of our custom CPMap kernel against traditional RBF, poly, and linear kernels using Parkinson’s disease dataset. We employed statistical significance testing to confirm the superiority of CPMap in terms of accuracy and the MCC.
@@
This statistical analysis and visualization confirm the significant advantage of the CPMap kernel over traditional kernels in the context of Parkinson’s disease data, reinforcing the potential for quantum machine learning to achieve empirical improvement in predictive tasks.
@@ II.4.3 Test with PCA reduced dataset on ibm_torino
In this case, we deployed our CPMap kernel circuit on IBM’s ibm_torino, a 133-qubit superconducting quantum processor belonging to the Heron family. Heron-class devices feature improvements in qubit coherence and native two-qubit gate implementations. At the time of execution, the backend-reported median coherence times were approximately $T_{1}\approx 1.83\times 10^{-4}\,\mathrm{s}$ and $T_{2}\approx 1.41\times 10^{-4}\,\mathrm{s}$.
Due to limited hardware access, we performed this experiment using a PCA-reduced representation with $d=7$ features, encoded on 4 qubits using CPMap. The overlap circuits required for the fidelity kernel were executed on hardware using IBM Runtime’s Sampler primitive with otherwise default runtime settings. We used 1024 shots per circuit and did not apply any error mitigation or correction techniques for this result. We did not manually select a qubit layout; instead, we allowed the IBM Runtime compilation workflow to choose the physical qubits and initial layout automatically. All circuits were transpiled with optimization level 3.
@@
The quantum kernel matrix was computed in a single run (i.e., without recomputing entries across training and test evaluation), and the training and test kernel submatrices were extracted from this precomputed kernel. We allocated 80 samples for training and 20 samples for testing.
@@
The observed performance gap between the statevector simulation and the real-device execution can be attributed to decoherence time, gate and measurement errors in the compiled overlap circuits. Nevertheless, the successful deployment of CPMap on ibm_torino confirms its feasibility on today’s NISQ devices and underscores the promise of quantum kernels in practical machine learning tasks.
Figure 11: Comparison of CPMap based kernel on the curated Parkinson dataset with PCA-reduced input ($d=7$), executed on ibm_torino (Sampler, 1024 shots, optimization level 3; no error mitigation). The simulated CPMap-based Kernel achieves the highest performance, followed by RBF and quantum hardware (ibm_torino).
II.5 Parameter sensitivity and kernel geometry
CPMap contains a small set of fixed design parameters that determine how pairs of qubits interact inside the feature map. These parameters are not trained, so it is important to check whether the method is stable and whether parameter variations induce meaningful (rather than arbitrary) changes in the resulting kernel. To do this, we performed a random parameter search on a 400-sample subset of the Breast Cancer dataset: we randomly sampled CPMap angle settings (varying both the C and P-block parameters), constructed the corresponding kernel matrices, and evaluated the same SVM pipeline using the mean performance over 50 stratified train//test splits. For each sampled setting, we also computed kernel–label alignment, a standard scalar summary of how well the kernel similarity matrix matches the class structure (higher alignment means that samples from the same class tend to be more similar under the kernel than samples from different classes). As shown in Fig. 12, parameter settings with higher alignment consistently yield higher classification performance (correlation $\approx$ 0.57). This indicates that the CPMap parameters control the geometry of the induced kernel in a systematic way, rather than acting as arbitrary constants, and it provides a practical diagnostic for selecting reasonable parameter regimes. Another point to note is that the performance varies substantially across sampled settings (a wide spread in MCC), demonstrating that the CP parameters are not innocuous constants; rather, they can materially influence the quality of the kernel and downstream accuracy. Taken together, these results support two practical conclusions: (i) CPMap is not overly brittle—many parameterizations remain competitive—but (ii) the choice of angles does matter, and alignment provides a simple diagnostic for identifying promising regimes without introducing trainable parameters. Additional experimental details and kernel diagnostics are provided in Supplementary Section G.
Figure 12: Kernel geometry vs performance under random parameter search on Breast Cancer ($n=400$, $d=12$). Each point is one parameter setting evaluated over 50 stratified splits. The x-axis is kernel–label alignment (higher means the kernel similarities better match class structure) and the y-axis is mean MCC. A strong positive trend (corr $\approx 0.57$) shows that CPMap parameters systematically shape kernel geometry in a way that predicts performance. The wide dispersion in MCC further shows that the parameters materially affect results. The solid line is a least-squares fit.
II.6 Issue of Exponential Concentration:
@@ III Discussion and Conclusion
Our proposed feature map paves the way for the application of quantum machine learning algorithms to more complex and higher-dimensional data, even with limited qubit resources. The CPMap’s remarkable efficiency in qubit utilization, evident through its requirement of approximately $\frac{N}{2}$ qubits to represent $N$ features, marks a substantial improvement over the ZZFeatureMap. This efficiency not only eases the computational load but also enhances the scalability of quantum models. In our comparative analysis with classical kernels such as the RBF, linear, and polynomial kernels, the CPMap sometimes demonstrated superior performance. Our findings show that, for the studied datasets and preprocessing choices, CPMap can achieve competitive (and sometimes higher) predictive performance than standard classical kernels, while using substantially fewer quantum resources than common Pauli feature maps.
@@
Note on classical simulability.
CPMap is intentionally shallow and locally structured to improve NISQ feasibility. The circuit structure underlying CPMap resembles QCNN-style architectures. Such architectural choices can also place a model in regimes where classical approximation methods remain effective. Recent work by Bermejo [50] argues that QCNNs can be effectively classically simulable in practically relevant regimes, most notably, when their action is restricted to information accessible via low-bodyness observables and when the benchmark instances are “locally-easy,” so that a classical algorithm equipped with (Pauli) classical shadows can reproduce the relevant behaviour.
In contrast, our method uses CPMap as a feature map, for kernel method, where training and evaluation require estimating state overlaps of the form $|\langle 0|U(y)^{\dagger}U(x)|0\rangle|^{2}$ across many pairs. Whether such overlap estimation is efficiently approximable classically can depend strongly on the circuit depth, entanglement growth, and the input distribution (e.g., the tensor-network bond dimension required for accurate contraction). We therefore do not make a general simulability or hardness claim here; rather, we position CPMap as a resource-efficient embedding and treat questions of classical simulability as regime-dependent.
The implications of this work extend beyond the specific benchmarks considered here. By reducing qubit requirements and limiting the growth of two-qubit entangling operations, CPMap makes kernel-based quantum learning experiments more accessible at higher feature counts, where many standard feature maps quickly become resource-intensive. This motivates several directions for future work, particularly in applying quantum algorithms to more complex and diverse datasets with applications in health sciences to material discovery. Further exploration into its realization on existing noisy quantum computing platforms and future small-scale fault-tolerant devices, as opposed to simulations, could offer deeper insights into its practical utility and performance in real-world applications.
Rather than positioning CPMap as a universal replacement for Pauli feature maps, we view it as a resource-efficient alternative that can be advantageous when qubit count and entangling-gate budgets are the dominant constraints. Our numerical benchmarks and small-scale hardware demonstrations indicate that CPMap can achieve competitive performance while using fewer qubits and fewer entangling gates (CNOTs) than standard Pauli-style constructions at comparable feature counts. These results suggest CPMap as a useful candidate for near-term kernel workflows and for systematic benchmarking against strong classical baselines. Clarifying the regimes where CPMap provides consistent benefits—across larger datasets, wider classes of quantum embeddings, and more realistic noise models—remains an important next step.